OpenStudy (ltrout):

[Fan + Medal] Find all of the zeros of the equation: -3x^4 + 27x^2 + 1200 = 0. Please don't just give the answer, work it out!

OpenStudy (anonymous):

need help

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

put \(u=x^2\) and solve \[-3u^2+27u+1200=0\] for \(u\)

OpenStudy (priyar):

yeah..now its simply a quad eq..

OpenStudy (misty1212):

probably easiest to divide by \(-3\) next, get \[u^2-9u-400=0\]

OpenStudy (anonymous):

http://www.bing.com/search?q=-3x%5E4+%2B+27x%5E2+%2B+1200+%3D+0.&form=IE11TR&src=IE11TR&pc=DCJB here is another way I found out so you can check and it is very precise

OpenStudy (misty1212):

then if you are lucky it factors

OpenStudy (ltrout):

@misty1212 Why do we change x^2 to u?

OpenStudy (misty1212):

because you can always solve a quadratic equation

OpenStudy (misty1212):

then change back at the end you don't really have to do that, if you know how to factor \[ -3x^4 + 27x^2 + 1200 \]

OpenStudy (misty1212):

it is just a gimmick to see it easily

OpenStudy (ltrout):

Okay @misty1212 , I got it:). Could you help on one more?

OpenStudy (misty1212):

i gotta run to class now good luck